Centre of circle is the limit of point of intersection of the lines 3x + 5y =1 and (2+c)x + 5(c^2)y =1 as c tends to 1. If it passes through (2,0) and its radius is p\q then p+q equals to (p and q are coprime)

Grade:12

1 Answers

Meet
6 years ago
If we put the value of x from 1st equation to second equation then we get x= 2/5 and y= -1/25 and this will lead to centre of circle.